OpenStudy (anonymous):

what is the square root of 6/7 ( rationalize the denominator and simplify)

OpenStudy (hunus):

\[\sqrt{\frac{6}{7}}=\sqrt{\frac{6}{7}}*\sqrt{\frac{7}{7}}=\frac{\sqrt{42}}{7}=\frac{\sqrt{2*7*3}}{7}=\frac{\sqrt{42}}{7}\]